# Suhyeon Lee    Dieu-Huyen Nguyen    Donghwan Lee

### v1

@@ 1 Introduction
Blockchains offer a decentralized, verifiable execution environment. Because many nodes must replicate state, on-chain computation is costly. To scale, systems shift intensive work off-chain and verify correctness only when disputes arise through challenge protocols, including Truebit-style verification, fraud proofs in optimistic rollups, and data-availability audits. These mechanisms keep the common case cheap while retaining a safety valve: if at least one honest party challenges, fraud can be surfaced and a slashed deposit funds the payout. It seems very easy to satisfy. However, this impression is not enough to construct a blockchain protocol for users considering its nature. For example, if a dishonest proposer is disputed by a challenger account which is actually the proposer’s Sybil account, the dishonest proposer will be able to avoid any loss. In this paper, we raise two questions. First, are honest challengers are incentivized in the blockchain environments? The first question is to guarantee at least one honest validator exists. Second, do dishonest proposers who post wrong values get proper economic damage that the system intends? The second question is to guarantee dishonest proposers do not appear.
@@ 3 Model and Design Goals
In this section, we formalize a minimal model for challenge-based rewards: players, costs with an upper bound, deposits, ordering modes, and winner multiplicity. We state two design goals used in the following sections.

### v2

@@ 1 Introduction
Blockchains offer a decentralized, verifiable execution environment. Because many nodes must replicate state, on-chain computation is costly. To scale, systems shift intensive work off-chain and verify correctness only when disputes arise through challenge protocols, including Truebit-style verification, fraud proofs in optimistic rollups, and data-availability audits. These mechanisms keep the common case cheap while retaining a safety valve: if at least one honest party challenges, fraud can be surfaced and a slashed deposit funds the payout. It seems very easy to satisfy. However, this impression is not enough to construct a blockchain protocol for users considering its nature as verifier’s dilemma [8]. Also, even if an honest verifier is not lazy, it does not guarantee that the verification will be compensated enough. For example, if a dishonest proposer is disputed by a challenger account which is actually the proposer’s Sybil account, the dishonest proposer will be able to avoid any loss. In this paper, we raise two questions. First, are honest challengers incentivized in the blockchain environments? The first question is to guarantee at least one honest validator exists. Second, do dishonest proposers who post wrong values get proper economic damage that the system intends? The second question is to guarantee dishonest proposers do not appear.
@@ 3 Model and Design Goals
In this section, we formalize a minimal model for challenge-based rewards: players, costs with an upper bound, deposits, ordering modes, and winner multiplicity. This paper follows a mechanism design approach, which can be conceptualized as the inverse of traditional game theory analysis, in which the objective is designing game rules to achieve desired outcomes instead of analyzing all possible outcomes from existing, fixed rules. We state two design goals used in the following sections.
@@ 3.2 Reward pool, winners, and splits
3.4 Key Parameters and Payoff Matrix
We consider a game between an Honest Challenger (HC) and the Adversarial Coalition (AC). The HC always challenges a potentially fraudulent transaction, while the AC decides whether to Commit Fraud. We assume that once an honest challenger initiates a dispute, the fraud is always detected. The loss of the adversarial coalition is determined as $D_{p}-\phi\alpha D_{p}$ = $(1-\phi\alpha)D_{p}$. This loss is proportional to the size of coalition $\phi$ and the reward fraction $\alpha$ of proposer deposit if fraud proof success. A summary of notations is provided in Table 1, and the detailed payoff matrix for the single-winner game is presented in Table 2.
Table 1: Description of Parameters
Symbol Parameter Description $c_{i}^{init}$ Challenger Initial Cost The initial cost for an honest challenger to verify and submit a challenge. $c_{i}^{proc}$ Challenger Processing Cost The processing cost for an honest challenger. $c_{i}$ Total Challenger Cost The total cost for an honest challenger, including initial and processing costs. $D_{p}$ Proposer’s Deposit The deposit made by the proposer. $\alpha$ Reward Fraction The fraction of the proposer’s deposit ($D_{p}$) distributed as a reward if fraud proof succeeds. $\phi$ Colluding Fraction The fraction of colluding challengers within the adversarial coalition (equivalent to $A/N$). $\eta$ Min. Deterrence Level The minimum deterrence level required to prevent fraud. $f$ Priority Fee The priority fee a validator pays in a U-b or U-P setup. $f^{*}$ Market-Clearing Fee The market-clearing priority fee.
Table 2: Payoff Matrix for Single-Winner Game
Challenger Action AC: Commit Fraud AC: Not Commit Fraud
HC: Challenge (U-P) $\alpha D_{p}-c_{i}-f^{*};-(1-\phi\alpha)D_{p}$ $c_{i}-f^{*};0$
HC: Challenge (U-b) $\alpha D_{p}-c_{i}-f^{*};-(1-\phi\alpha)D_{p}$ $c_{i}-f^{*};0$
HC: Challenge (Fair) $\alpha D_{p}-c_{i};-(1-\phi\alpha)D_{p}$ $c_{i};0$
Under the multi-winner setting, the economic value of transaction ordering is significantly diminished. Therefore, the ordering fee is nearly zero with all the order fairness environments (U-P, U-b, and F). Each of the $N$ included winners is eligible for $\frac{1}{N}$ of the reward. Table 3 illustrates the payoff matrix under this multi-winner setting.
Table 3: Payoff Matrix for Multi-Winner Game
Challenger Action AC: Commit Fraud AC: Not Commit Fraud
HC: Challenge $\frac{1}{N}\alpha D_{p}-c_{i};-(1-\phi\alpha)D_{p}$ $c_{i};0$